# Sparse Representations of Dynamical Networks:

### v4

@@ 1 Introduction
With the growing trend of modern technical applications becoming more interconnected and spatially widespread, distributed process control has become an area of intense research in the specialized literature catering to dynamical systems. However, with its convexity refuted and its generic intractability revealed (see, for example, [3]), recent efforts in the area of distributed control for Linear and Time-Invariant (LTI) systems have shifted from the classical formulation of imposing sparsity upon the Transfer Function Matrix (TFM) of the controller [32, 33, 22, 26] towards the implementation of decentralized state-space-based control laws.
@@ 1.1 Context
Among the first attempts to formalize a state-space-oriented approach to distributed control was the work from [17], in which structured realizations are employed in order to obtain control laws that inherit the sparsity structure of the controlled network. A key concept pioneered in [17], which turns out to be crucial for the results presented in this manuscript, is the fact that the distributed state-space implementation of the sought-after control laws must possess certain structural properties, such as stabilizability and detectability, in order to certify closed-loop stability.
Yet, arguably, the most celebrated results in this new paradigm of distributed control came with the advent of the System-Level Synthesis (SLS) design framework from [35]. Based upon the classical work in [1], the SLS approach produces decentralized control laws for discrete-time systems, by employing an affine parametrization of all the achievable closed-loop maps. Additionally, the work in [13] has shown that, in the SLS framework, it is possible to decouple controller synthesis from closed-loop design. Closely following the developments in [35], the body of work presented in [10] aims to provide a continuous-time parallel to the SLS framework, with important contributions to the vehicular platoon problem, which was also tackled in [25]. Notably, the approaches from [10] and [25] are linked via the implicit use of coprime factorizations, and an underlying dependence upon the Youla parametrization.
In the same vein as the techniques from [10] and [25], a new design paradigm, based upon coprime factorizations and heavily inspired by the work from [37], was recently formalized in [24], with the aim of proposing a unitary continuous- and discrete-time approach to optimal distributed control. The theoretical developments presented in [24] are also supported by the convex relaxation-based procedures from [27], which are part of a larger class of design methods currently gaining traction in the distributed control literature (see [5, 14, 16, 34, 23]). Fortunately, the intrinsic connections between the SLS-focused methods and their Youla-based counterparts has been elucidated in [29] and [39], while a system-theoretical middle ground between these two approaches was proposed in [7] and [38], via the Input-Output Parametrization.
@@ 1.2 Motivation
Although the various techniques mentioned in the previous subsection clearly show that significant progress has been made in moving away from the traditional sparse TFM-based approach to distributed control, a number of important limitations still remain unaddressed in literature. While certain procedures, such as the ones in [17], build upon highly structured realizations for the investigated networks, obtaining these exceptional representations becomes somewhat unwieldy for large-scale systems. Except for those networks endowed with particularly sparse or with recursive inter-nodal connections (see the numerical example in [27]), the effort needed to compose the interconnected plant’s model out of its nodes’ realizations does not scale gracefully with respect to the network’s complexity.
Similar to the results from [17], other methods proposed in literature focus exclusively on networks modeled in either continuous- or discrete-time, with [10] and [35] serving as respective examples. Oftentimes, this exclusivity is owed to the manner in which stability guarantees are offered for the proposed control schemes. We point out that this particular limitation, which will be further commented upon in the sequel, has been lifted following the introduction of the design framework from [24]. In doing so, however, the control laws from [24] take on a distinct input-output interpretation, regardless of their distributed and stabilizing state-space-based implementations, as in Section III.C in [24]. Thus, the cost paid for this increased flexibility is a degree of opaqueness with respect to network state dynamics, at the implementation level.
@@ 1.3 Contribution
The aim of this paper is to propose a factorized representation, inspired by the system response TFMs from [35], which is suitable to both network analysis and distributed controller synthesis. This representation, dubbed a System Response-Type Realization (SRTR), captures both the dynamics and the structure of the network, regardless of whether it is modeled as either a continuous- or a discrete-time system. This factorization offers a notably keener insight into the state dynamics of the plant than the Network Realization Function (NRF) representation from [24], while ensuring the same guarantees of closed-loop stability as NRF-based control laws, when implementing controllers via SRTR-based feedback schemes.
In developing the theory associated with this new representation, we implicitly investigate the subtle connections between the classical left coprime factorizations discussed in [31], the system response TFMs from [35] and the NRF pairs introduced in [24]. Additionally, we propose numerically tractable means of shifting between the aforementioned types of representations, all while placing a particular emphasis on the preservation of three key properties: stability, coprimeness and sparsity structure. Finally, we show how the first property allows us to elegantly tackle the problem of strong stabilization in the decentralized setting, how the second property implicitly allows for the commutation between representations, and how the third property directly promotes the implementation of distributed control laws.
@@ 1.4 Outline of the Paper
In the second section, we include a series of preliminary notions which are standard to control theory literature. The third section of the paper contains a brief outline of the theoretical concepts behind various representations for LTI systems. In the fourth section, we show that, while the NRF representation of a given LTI system is not guaranteed to be coprime, this property can be ensured for a closely related representation associated with our system, heavily inspired by those in [35]. We also provide state-space formulas for the class of all such coprime representations, related to a particular type of realization for our LTI system.
The fourth section also contains the main results of the paper: an explanation of the intrinsic connections between NRF pairs, the coprime representations inspired by [35], and the classical left coprime factorizations associated with a given TFM, along with the means of employing them to obtain distributed and stabilizing control laws. The fifth section showcases the direct application of these results through a numerical example. Finally, the sixth section contains a number of concluding remarks.
@@ 2.1 Notation
[…]
@@ 2.2 Realization Theory
A pair $(A,B)$ or a realization (1a)-(1b) for which $\begin{bmatrix}A-\lambda I_{n}&B\end{bmatrix}$ has full row rank at $\lambda\in\mathbb{C}$ is called controllable at $\lambda$. Similarly, a pair $(C,A)$ or a realization (1a)-(1b) that is observable at $\lambda$ is one for which the pair $(A^{\top},C^{\top})$ is controllable at $\lambda$. A controllable/observable matrix pair or realization (1a)-(1b) is one that is controllable/observable at any $\lambda\in\mathbb{C}$. A realization whose order is the smallest out of all others for the same $\mathbf{G}(\lambda)$ is called minimal (see Section 6.5.1 of [12]). Any two minimal realizations $(A,B,C,D)$ and $(\widehat{A},\widehat{B},\widehat{C},\widehat{D})$ of $\mathbf{G}(\lambda)$ having order $\nu$ are related via a state equivalence transformation $x=T{\widehat{x}}$, with $T$ invertible in $\mathbb{R}^{\nu\times\nu}$, given by
(2) $\begin{array}[]{llll}A=T\widehat{A}T^{-1},&B=T\widehat{B},&C=\widehat{C}T^{-1},&D=\widehat{D}.\end{array}$
@@
Let $\mathbb{S}$ denote the domain of stability, which is either the open left-half plane, i.e., $\left\{\lambda\in\mathbb{C}\ :\ \text{Re}(\lambda)<0\right\}$ in the case of continuous-time systems or the open unit disk, i.e., $\left\{\lambda\in\mathbb{C}\ :\ |\lambda|<1\right\}$ for discrete-time ones. A stabilizable realization (1a)-(1b) or pair $(A,B)$ is one that is controllable at any $\𝕊\lambda\in\mathbb{C}\backslash\mathbb{S}$. By duality, a detectable realization (1a)-(1b) or pair $(C,A)$ is one that is observable at any $\𝕊\lambda\in\mathbb{C}\backslash\mathbb{S}$.
@@ 2.3 Rational Matrix Theory
The matrix polynomial $A-\lambda E$ is called a pencil. If it is square and $\det(A-\lambda E)\not\equiv 0$, we refer to it as regular (see chapter 12 of [8] for more details). Note also that a pencil is merely a polynomial TFM. The finite poles and zeros of any TFM $\mathbf{G}(\lambda)$ are rigorously defined in terms of the Smith-McMillan Form (see Section 6.5.2 of [12]), whereas the infinite poles and zeros of $\mathbf{G}(\lambda)$ are those of $\mathbf{G}\left(\frac{1}{\lambda}\right)$ at $\lambda=0$. Any realization (1a)-(1b) that is both controllable and observable is also minimal (see [36]) and the poles of its TFM coincide with the eigenvalues of the regular pencil $A-\lambda I_{n}$, known as the realization’s pole pencil.
[…]
Two TFMs $\mathbf{N}(\lambda)\in\mathbb{R}(\lambda)^{p\times m}$ and $\mathbf{M}(\lambda)\in\mathbb{R}(\lambda)^{p\times p}$, with $\mathbf{M}(\lambda)$ having full normal rank, that satisfy $\mathbf{G}(\lambda)=\mathbf{M}^{-1}(\lambda)\mathbf{N}(\lambda)$ are said to be a left factorization of $\mathbf{G}(\lambda)$. When, additionally, $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}$ has full normal rank (this is always ensured by $\mathbf{M}(\lambda)$ being invertible) along with no zeros in $\mathbb{C}_{\infty}$ (see also [20]), we say that $\mathbf{M}(\lambda)$ and $\mathbf{N}(\lambda)$ constitute a fully left coprime factorization (FLCF) of $\mathbf{G}(\lambda)$ (see [31, 20, 21] for equivalent characterizations). When $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}$ has full row normal rank along with no poles or zeros located outside of $\mathbb{S}$, the pair is called an LCF over $\mathbb{S}$.
@@ 3 Representations for LTI Networks
This section contains a discussion based upon various representations for networks of LTI systems, with an emphasis on those tentatively prototyped in [28] and now discussed in a broader context.
@@ 3.1 State-Space Realizations
Note that Assumption 3.1 is in no way restrictive from a practical standpoint, where oftentimes $n\gg p$. The lack of full row rank for the matrix $\widetilde{C}$ implies that some of the sensor measurements are redundant and can be amalgamated via linear combinations to produce a new, full row rank $C$-matrix, all while preserving the property of observability (see Section 3.4 of [36] for the underlying connection with reduced-order observers). Notice also that when $n=p$, the matrix $\widetilde{C}$ is invertible and the problem can be reduced to the trivial, full state measurement case, in which the main representations discussed in the sequel take on the classical form given in (3a).
@@
[…]
@@
$\lambda\mathbf{Y}(\lambda)=\big{(}A_{11}+A_{12}(\lambda I_{n-p}-A_{22})^{-1}A_{21}\big{)}\mathbf{Y}(\lambda)+\big{(}B_{1}+A_{12}(\lambda I_{n-p}-A_{22})^{-1}B_{2}\big{)}\mathbf{U}(\lambda),$
@@ 3.2 Network Realization Functions
We now formally introduce the notion of NRF pair, discussed at length in [24].
@@
Note that the outputs of the network’s nodes need not be scalar and that constraining only the diagonal entries of $\mathbf{\Phi}(\lambda)$ represents the most generic case. Indeed, when implementing controllers via their NRF pairs, it may be convenient for certain nodes to compute more than one command signal, with the latter representing our network’s outputs. Therefore, it is natural to impose identically zero elements on the block-diagonal of $\mathbf{\Phi}(\lambda)$, in order to avoid the redundant use of local information. This merely amounts to an additional sparsity constraint on the entries of $\mathbf{\Phi}(\lambda)$ and we point out that the results from [24] are not dependent upon the structure of the aforementioned TFM, except for its identically zero diagonal. Moreover, as explained in Remark III.8 of [24], the block-row implementations which arise naturally from imposing identically zero blocks on the block-diagonal of $\mathbf{\Phi}(\lambda)$ retain all the desirable properties of the row implementations provided in the results of Section III from [24].
@@
Another benefit of the these NRF representations consists in the preservation of the network’s adjacency graph, both with respect to the input-output relationships that characterize it, as modeled by $\mathbf{\Gamma}(\lambda)$, and also of the intrinsic interconnections of its manifest variables, as encoded through the structure of $\mathbf{\Phi}(\lambda)$. These graph-oriented representations have received significant attention in the context of Kron reduction for electrical networks (see [2, 4, 6] and the references therein).
@@ 3.3 An Example of NRF Modelling
(11) $\small\begin{bmatrix}\mathbf{U}_{1}(\lambda)\\ \mathbf{U}_{2}(\lambda)\\ \mathbf{U}_{3}(\lambda)\end{bmatrix}=\small\begin{bmatrix}O&O&\mathbf{\Phi}_{1}(\lambda)\\ \mathbf{\Phi}_{2}(\lambda)&O&O\\ O&\mathbf{\Phi}_{3}(\lambda)&O\end{bmatrix}\normalsize\small\begin{bmatrix}\mathbf{U}_{1}(\lambda)\\ \mathbf{U}_{2}(\lambda)\\ \mathbf{U}_{3}(\lambda)\end{bmatrix}+\small\begin{bmatrix}\mathbf{\Gamma}_{1}(\lambda)&O&O\\ O&\mathbf{\Gamma}_{2}(\lambda)&O\\ O&O&\mathbf{\Gamma}_{3}(\lambda)\end{bmatrix}\normalsize\small\begin{bmatrix}\mathbf{Z}_{1}(\lambda)\\ \mathbf{Z}_{2}(\lambda)\\ \mathbf{Z}_{3}(\lambda)\end{bmatrix},\normalsize$
@@
(12) $\mathbf{\Phi}(\lambda):=\small\begin{bmatrix}O&O&\mathbf{\Phi}_{1}(\lambda)\\ \mathbf{\Phi}_{2}(\lambda)&O&O\\ O&\mathbf{\Phi}_{3}(\lambda)&O\end{bmatrix},\quad\mathbf{\Gamma}(\lambda):=\small\begin{bmatrix}\mathbf{\Gamma}_{1}(\lambda)&O&O\\ O&\mathbf{\Gamma}_{2}(\lambda)&O\\ O&O&\mathbf{\Gamma}_{3}(\lambda)\end{bmatrix},\normalsize$
@@
Additionally, we wish to further illustrate how the NRF determines, via (10), the topology of the LTI network describing the given TFM $\mathbf{G}(\lambda)$. By considering $\mathbf{\Phi}_{13}(\lambda)$ identically zero, which would imply the “breaking” of the ring network from Fig. 1, the latter becomes a cascade connection and $\mathbf{G}(\lambda)$ can be seen as a “line” network. This is relevant in the motion control of vehicles moving in “platoon” formation [25].

### v7

@@ 1 Introduction
With the growing trend of modern technical applications becoming more interconnected and spatially widespread, distributed process control has become an area of intense research in the specialized literature catering to dynamical systems. However, with its convexity refuted and its generic intractability revealed (see, for example, [DateChow]), recent efforts in the area of distributed control for Linear and Time-Invariant (LTI) systems have shifted from the classical formulation of imposing sparsity upon the Transfer Function Matrix (TFM) of the controller [Voul1, Voul2, Rotko, Shah] towards the implementation of decentralized state-space-based control laws.
@@ 1.1 Context
Among the first attempts to formalize a state-space-oriented approach to distributed control was the work from [Realiz], in which structured realizations are employed in order to obtain control laws that inherit the sparsity structure of the controlled network. A key concept pioneered in [Realiz], which turns out to be crucial for the results presented in this manuscript, is the fact that the distributed state-space implementation of the sought-after control laws must possess certain structural properties, such as stabilizability and detectability, in order to certify closed-loop stability.
Yet, arguably, the most celebrated results in this new paradigm of distributed control came with the advent of the System-Level Synthesis (SLS) design framework from [SLS]. Based upon the classical work in [BBS], the SLS approach produces decentralized control laws for discrete-time systems, by employing an affine parametrization of all the achievable closed-loop maps. Additionally, the work in [Separ] has shown that, in the SLS framework, it is possible to decouple controller synthesis from closed-loop design. Closely following the developments in [SLS], the body of work presented in [Explicit] aims to provide a continuous-time parallel to the SLS framework, with important contributions to the vehicular platoon problem, which was also tackled in [plutonizare]. Notably, the approaches from [Explicit] and [plutonizare] are linked via the implicit use of coprime factorizations, and an underlying dependence upon the Youla parametrization.
In the same vein as the techniques from [Explicit] and [plutonizare], a new design paradigm, based upon coprime factorizations and heavily inspired by the work from [DSF], was recently formalized in [NRF], with the aim of proposing a unitary continuous- and discrete-time approach to optimal distributed control. The theoretical developments presented in [NRF] are also supported by the convex relaxation-based procedures from [aug_sparse], which are part of a larger class of design methods currently gaining traction in the distributed control literature (see [Lavaei, Jovanovic, reg, Sznaier, trian]). Fortunately, the intrinsic connections between the SLS-focused methods and their Youla-based counterparts has been elucidated in [Tseng] and [Luca3], while a system-theoretical middle ground between these two approaches was proposed in [Luca1] and [Luca2], via the Input-Output Parametrization.
@@ 1.2 Motivation
Although the various techniques mentioned in the previous subsection clearly show that significant progress has been made in moving away from the traditional sparse TFM-based approach to distributed control, a number of important limitations still remain unaddressed in literature. While certain procedures, such as the ones in [Realiz], build upon highly structured realizations for the investigated networks, obtaining these exceptional representations becomes somewhat unwieldy for large-scale systems. Except for those networks endowed with particularly sparse or with recursive inter-nodal connections (see the numerical example in [aug_sparse]), the effort needed to compose the interconnected plant’s model out of its nodes’ realizations does not scale gracefully with respect to the network’s complexity.
Similar to the results from [Realiz], other methods proposed in literature focus exclusively on networks modeled in either continuous- or discrete-time, with [Explicit] and [SLS] serving as respective examples. Oftentimes, this exclusivity is owed to the manner in which stability guarantees are offered for the proposed control schemes. We point out that this particular limitation, which will be further commented upon in the sequel, has been lifted following the introduction of the design framework from [NRF]. In doing so, however, the control laws from [NRF] take on a distinct input-output interpretation, regardless of their distributed and stabilizing state-space-based implementations, as in Section III.C in [NRF]. Thus, the cost paid for this increased flexibility is a degree of opaqueness with respect to network state dynamics, at the implementation level.
@@ 1.3 Contribution
The aim of this paper is to propose a factorized representation, inspired by the system response TFMs from [SLS], which is suitable to both network analysis and distributed controller synthesis. This representation, dubbed a System Response-Type Realization (SRTR), captures both the dynamics and the structure of the network, regardless of whether it is modeled as either a continuous- or a discrete-time system. This factorization offers a notably keener insight into the state dynamics of the plant than the Network Realization Function (NRF) representation from [NRF], while ensuring the same guarantees of closed-loop stability as NRF-based control laws, when implementing controllers via SRTR-based feedback schemes.
In developing the theory associated with this new representation, we implicitly investigate the subtle connections between the classical left coprime factorizations discussed in [V], the system response TFMs from [SLS] and the NRF pairs introduced in [NRF]. Additionally, we propose numerically tractable means of shifting between the aforementioned types of representations, all while placing a particular emphasis on the preservation of three key properties: stability, coprimeness and sparsity structure. Finally, we show how the first property allows us to elegantly tackle the problem of strong stabilization in the decentralized setting, how the second property implicitly allows for the commutation between representations, and how the third property directly promotes the implementation of distributed control laws.
@@ 1.4 Outline of the Paper
In the second section, we include a series of preliminary notions which are standard to control theory literature. The third section of the paper contains a brief outline of the theoretical concepts behind various representations for LTI systems. In the fourth section, we show that, while the NRF representation of a given LTI system is not guaranteed to be coprime, this property can be ensured for a closely related representation associated with our system, heavily inspired by those in [SLS]. We also provide state-space formulas for the class of all such coprime representations, related to a particular type of realization for our LTI system.
The fourth section also contains the main results of the paper: an explanation of the intrinsic connections between NRF pairs, the coprime representations inspired by [SLS], and the classical left coprime factorizations associated with a given TFM, along with the means of employing them to obtain distributed and stabilizing control laws. The fifth section showcases the direct application of these results through a numerical example. Finally, the sixth section contains a number of concluding remarks.
@@ 2.1 Notation
We denote by $\mathbb{R}$ the set of real numbers and by $\mathbb{N}$ the set of natural ones, while $\mathbb{C}$ stands for the complex plane and $\mathbb{C}_{\infty}:=\mathbb{C}\cup\{\infty\}$ denotes the extended complex plane. Let $\mathbb{R}^{p\times m}$ be the set of $p\times m$ real matrices, for whose elements we denote by $\|\cdot\|$ the spectral norm, i.e., the largest singular value. Let also $\mathbb{R}(\lambda)^{p\times m}$ be the set of $p\times m$ TFMs, which are matrices having real-rational functions as entries. Henceforth, we denote such TFMs through the use of boldface symbols.
@@ 2.2 Realization Theory
[…]
@@
Let $\mathbb{S}$ denote the domain of stability, which is either the open left-half plane, i.e., $\left\{\lambda\in\mathbb{C}\ :\ \text{Re}(\lambda)<0\right\}$ in the case of continuous-time systems or the open unit disk, i.e., $\left\{\lambda\in\mathbb{C}\ :\ |\lambda|<1\right\}$ for discrete-time ones. A stabilizable realization (1a)-(1b) or pair $(A,B)$ is one that is controllable at any $\lambda\in\mathbb{C}\backslash\mathbb{S}$. By duality, a detectable realization (1a)-(1b) or pair $(C,A)$ is one that is observable at any $\lambda\in\mathbb{C}\backslash\mathbb{S}$.
@@ 2.3 Rational Matrix Theory
The matrix polynomial $A-\lambda E$ is called a pencil. If it is square and $\det(A-\lambda E)\not\equiv 0$, we refer to it as regular (see chapter 12 of [gantmacher] for more details). Note also that a pencil is merely a polynomial TFM. The finite poles and zeros of any TFM $\mathbf{G}(\lambda)$ are rigorously defined in terms of the Smith-McMillan Form (see Section 6.5.2 of [Kai]), whereas the infinite poles and zeros of $\mathbf{G}(\lambda)$ are those of $\mathbf{G}\left(\frac{1}{\lambda}\right)$ at $\lambda=0$. Any realization (1a)-(1b) that is both controllable and observable is also minimal (see [Wonham]) and the poles of its TFM coincide with the eigenvalues of the regular pencil $A-\lambda I_{n}$, known as the realization’s pole pencil.
The McMillan degree of $\mathbf{G}(\lambda)$ is the sum of the degrees of its poles (see Section 6.5.3 of [Kai]), both finite and infinite. A TFM’s normal rank, which is its rank for almost all points in $\mathbb{C}_{\infty}$, is also rigorously defined via the TFM’s Smith-McMillan Form. If a TFM has no poles at $\{\infty\}$, it is called proper. Otherwise, it is called improper. A TFM $\mathbf{G}\in\mathbb{R}(\lambda)^{p\times m}$ for which $\lim_{\lambda\rightarrow\infty}\mathbf{G}(\lambda)=O$ is called strictly proper. Finally, a TFM is deemed stable if it has no poles located in $\mathbb{C}_{\infty}\backslash\mathbb{S}$.
Two TFMs $\mathbf{N}(\lambda)\in\mathbb{R}(\lambda)^{p\times m}$ and $\mathbf{M}(\lambda)\in\mathbb{R}(\lambda)^{p\times p}$, with $\mathbf{M}(\lambda)$ having full normal rank, that satisfy $\mathbf{G}(\lambda)=\mathbf{M}^{-1}(\lambda)\mathbf{N}(\lambda)$ are said to be a left factorization of $\mathbf{G}(\lambda)$. When, additionally, $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}$ has full normal rank (this is always ensured by $\mathbf{M}(\lambda)$ being invertible) along with no zeros in $\mathbb{C}_{\infty}$ (see also [SIMAX]), we say that $\mathbf{M}(\lambda)$ and $\mathbf{N}(\lambda)$ constitute a fully left coprime factorization (FLCF) of $\mathbf{G}(\lambda)$ (see [V, SIMAX, Rose70] for equivalent characterizations). When $\begin{bmatrix}\mathbf{M}(\lambda)&\mathbf{N}(\lambda)\end{bmatrix}$ has full row normal rank along with no poles or zeros located outside of $\mathbb{S}$, the pair is called an LCF over $\mathbb{S}$.
@@ 3 Representations for LTI Networks
This section contains a discussion based upon various representations for networks of LTI systems, with an emphasis on those tentatively prototyped in [CDC2013] and now discussed in a broader context.
@@ 3.1 State-Space Realizations
Note that Assumption 3.1 is in no way restrictive from a practical standpoint, where oftentimes $n\gg p$. The lack of full row rank for the matrix $\widetilde{C}$ implies that some of the sensor measurements are redundant and can be amalgamated via linear combinations to produce a new, full row rank $C$-matrix, all while preserving the property of observability (see Section 3.4 of [Wonham] for the underlying connection with reduced-order observers). Notice also that when $n=p$, the matrix $\widetilde{C}$ is invertible and the problem can be reduced to the trivial, full state measurement case, in which the main representations discussed in the sequel take on the classical form given in (3a).
@@
(4b) $y$ $=\begin{bmatrix}\hskip 5.69054ptI_{p}&\hskip 6.40186ptO\hskip 4.97922pt\end{bmatrix}\begin{bmatrix}y\\ q\end{bmatrix}$.
@@
$\hskip-5.69054pt\lambda\mathbf{Y}(\lambda)=\big(A_{11}+A_{12}(\lambda I_{n-p}-A_{22})^{-1}A_{21}\big)\mathbf{Y}(\lambda)+\big(B_{1}+A_{12}(\lambda I_{n-p}-A_{22})^{-1}B_{2}\big)\mathbf{U}(\lambda),$
@@ 3.2 Network Realization Functions
We now formally introduce the notion of NRF pair, discussed at length in [NRF].
@@
Note that the outputs of the network’s nodes need not be scalar and that constraining only the diagonal entries of $\mathbf{\Phi}(\lambda)$ represents the most generic case. Indeed, when implementing controllers via their NRF pairs, it may be convenient for certain nodes to compute more than one command signal, with the latter representing our network’s outputs. Therefore, it is natural to impose identically zero elements on the block-diagonal of $\mathbf{\Phi}(\lambda)$, in order to avoid the redundant use of local information. This merely amounts to an additional sparsity constraint on the entries of $\mathbf{\Phi}(\lambda)$ and we point out that the results from [NRF] are not dependent upon the structure of the aforementioned TFM, except for its identically zero diagonal. Moreover, as explained in Remark III.8 of [NRF], the block-row implementations which arise naturally from imposing identically zero blocks on the block-diagonal of $\mathbf{\Phi}(\lambda)$ retain all the desirable properties of the row implementations provided in the results of Section III from [NRF].
@@
Another benefit of the these NRF representations consists in the preservation of the network’s adjacency graph, both with respect to the input-output relationships that characterize it, as modeled by $\mathbf{\Gamma}(\lambda)$, and also of the intrinsic interconnections of its manifest variables, as encoded through the structure of $\mathbf{\Phi}(\lambda)$. These graph-oriented representations have received significant attention in the context of Kron reduction for electrical networks (see [Kron1, Kron3, Kron2] and the references therein).
@@ 3.3 An Example of NRF Modelling
[…]
@@
[…]
@@
Additionally, we wish to further illustrate how the NRF determines, via (10), the topology of the LTI network describing the given TFM $\mathbf{G}(\lambda)$. By considering $\mathbf{\Phi}_{13}(\lambda)$ identically zero, which would imply the “breaking” of the ring network from Fig. 1, the latter becomes a cascade connection and $\mathbf{G}(\lambda)$ can be seen as a “line” network. This is relevant in the motion control of vehicles moving in “platoon” formation [plutonizare].